Turner Container Company is suffering declining sales of its principal product, nonbiodegradeable plastic cartons. The president, Robert Griffin, instructs his controller, Alexis Landrum, to lengthen asset lives to reduce depreciation expense. A processing line of automated plastic extruding equipment, purchased for $3.5 million in January 2019, was originally estimated to have a useful life of 8 years and a salvage value of $300,000. Depreciation has been recorded for 2 years on that basis. Robert wants the estimated life changed to 12 years total, and the straight-line method continued. Alexis is hesitant to make the change, believing it is unethical to increase net income in this manner. Robert says, “Hey, the life is only an estimate, and I've heard that our competition uses a 12-year life on their production equipment.” Instructions (a) Who are the stakeholders in this situation? (b) Is the change in asset life unethical, or is it simply a good business practice by an astute president? (c) What is the effect of Robert Griffin's proposed change on income before taxes in the year of change? if you have a source that i can rely on it will be great, also try to show the change with journalizing
Answer #1

a. Stakeholders: stakeholders means a group of persons or organizations that has interest in the organization

Examples:customers, Investors, directors, suppliers and vendors.

Here in the given case, the stakeholders are customers as the corporation is mostly based on the product based organization

B. The life of the asset is an estimate. The useful life of asset can be changed only when the estimate is reasonable and based on the industry and market conditions.

However, in the given case the sole purpose of revising the useful life of asset is to increase net income. Hence it is unethical.. The organization should disclose the reasons for changing the estimate.

Depreciation expense for previous period =( 3,500,000 -300, 0000) /8 = 400,000

Cumulative depreciation for 2 years = 800,000

The depreciation expense due to change in estimate

(Asset value - cumulative depreciation - salvage value) /remaining life of asset

= (3,500,0000 - 800,000 - 300,000) /10

= 240,000

Due to this change the net income of the organisation will increase by 160,0000 (400, 000 - 240,000)

There will be no retrospective effect in case of change in the accounting estimate.

Before change:

Depreciation 400,0000

Accumulated Depreciation 400,0000

After change:

Depreciation 240,0000

Accumulated depreciation 240,000